# An Interpretable Generalization Mechanism for Accurately Detecting Anomaly and Identifying Networking Intrusion Techniques

### v1

@@ 2.2 Why Should We Need an Interpretable-generalization IDS?
Ref.
Datasets
Features
Instances (Normal, Attack)
[3]
NSL-KDD
41 (All)
125973 (53%, 47%)
UNSW-NB15
47 (All)
118000 (47%, 53%)
[9]
NSL-KDD
12 from 41
148517 (52%, 48%)
[17]
UNSW-NB15
15 (5-19) from 41
First 200,000 (unknown)
[18]
UNSW-NB15
3 (18, 23, 26) from 47
2540044 (unknown)
[19]
UNSW-NB15
20 from 47
5576 (45%, 55%)
[20]
UNSW-NB15
8 from 47
32755 (50%, 50%)
[21]
UNSW-NB15
7 from 47
20000 (50%, 50%)
[22]
NSL-KDD
41 (All)
47600 (48%, 52%)
UNSW-NB15
47 (All)
257632 (36%, 64%)
[23]
UKM-IDS20
ARP poisoning: 2 from 48, DoS: 42 from 48, Scans: 40 from 48, Exploits: 40 from 48
Total: 12887 (69%, 31%); ARP poisoning: 8909, 592; DoS: 8909, 1742; Scans: 8909, 597; Exploits: 8909, 1047
@@ 1. Step 1:
[…]
@@ 2. Step 2:
[…]
@@ 4.2 Data Preprocessing
Dataset
Attacks: (Type, Amounts)
NSL-KDD
(DoS, 5388), (Probe, 1452), (R2L, 669), (U2R, 52)
UNSW-NB15
(Analysis, 500), (Backdoors, 500), (DoS, 500), (Exploits, 500), (Fuzzers, 500), (Generic, 500), (Reconnaissance, 500), (Shellcode, 500), (Worms, 174)
UKM-IDS20
(ARP poisoning, 592), (DoS, 1742) (Scans, 597), (Exploits, 1047)

### v2

@@ 2.2 Why Should We Need an Interpretable-generalization IDS?
Ref. Datasets Features Instances (Normal, Attack)
[3] NSL-KDD 41 (All) 125973 (53%, 47%)
UNSW-NB15 47 (All) 118000 (47%, 53%)
[9] NSL-KDD 12 from 41 148517 (52%, 48%)
[17] UNSW-NB15 15 (5-19) from 41 First 200,000 (unknown)
[18] UNSW-NB15 3 (18, 23, 26) from 47 2540044 (unknown)
[19] UNSW-NB15 20 from 47 5576 (45%, 55%)
[20] UNSW-NB15 8 from 47 32755 (50%, 50%)
[21] UNSW-NB15 7 from 47 20000 (50%, 50%)
[22] NSL-KDD 41 (All) 47600 (48%, 52%)
UNSW-NB15 47 (All) 257632 (36%, 64%)
[23] UKM-IDS20 ARP poisoning: 2 from 48, DoS: 42 from 48, Scans: 40 from 48, Exploits: 40 from 48 Total: 12887 (69%, 31%); ARP poisoning: 8909, 592; DoS: 8909, 1742; Scans: 8909, 597; Exploits: 8909, 1047
@@ 1. Step 1:
    For each test instance $T_{i}$, we calculate the normal score based on the occurrence and length of any CNP that is a subset of $T_{i}$. The normal-score $NS_{T_{i}}$ is determined as Equation (6), where CNPν is a CNP that is a subset of $T_{i}$.
@@ 2. Step 2:
    Similarly, an anomaly score $AS_{T_{i}}$ is computed for each test instance if a CAP is its subset, which is determined by Equation (7) and CAPω is a CAP that is a subset of $T_{i}$.
@@ 4.2 Data Preprocessing
Dataset Attacks: (Type, Amounts)
NSL-KDD (DoS, 5388), (Probe, 1452), (R2L, 669), (U2R, 52)
UNSW-NB15 (Analysis, 500), (Backdoors, 500), (DoS, 500), (Exploits, 500), (Fuzzers, 500), (Generic, 500), (Reconnaissance, 500), (Shellcode, 500), (Worms, 174)
UKM-IDS20 (ARP poisoning, 592), (DoS, 1742) (Scans, 597), (Exploits, 1047)